An enterprising student has set up an internship clearinghouse for business students. Each

student who uses the service fi lls out a form and lists up to 10 companies that he or she

would like to have contacted. The clearinghouse has a choice of two methods to use for processing

the forms. The traditional method requires about 20 minutes to review the form and

arrange the information in the proper order for processing. Once this setup is done, it takes

only two minutes per company requested to complete the processing. The other alternative

uses an optical scan/retrieve system, which takes only a minute to prepare but requires fi ve

minutes per company for completing the processing. If it costs about the same amount per

minute for processing with either of the two methods, when should each be used?
Answer #1

Time for processing 1 form:

Method 1

Form Processing = 20 minutes

Each Company processing = 2 minutes

Total Time for processing = 20 + 2*10 = 40

Method 2

Form Processing = 1 minutes

Each Company processing = 5 minutes

Total Time for processing = 1 + 5*10 = 51

Hence method 1 is better as it is faster